### v2

@@ 1 Introduction
Figure 1: Procedural Mistake Detection involves identifying errors within a procedural video. Each procedure is composed of different steps that should be executed with a certain order. The aim is to develop a method capable of analyzing a video and determining whether each frame contains a mistake, e.g., a step that is not executed in the correct order. Addressing this task is particularly useful when applied to wearable devices, as they allow for direct feedback to the person performing the task. In the Figure, TI-PREGO, depicted by a gear icon, takes as input the video sequence from time 00 to $t-1$, classifying the current frame $t$ as either correct (green) or a mistake (red). This process continues frame by frame until a mistake is detected.
@@
Advanced mistake detection models will become essential to ensure precision, safety, and efficiency in procedural applications. This demand has recently resulted in an increasing number of datasets [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] and methodologies designed to advance procedure learning [13, 4, 14, 15, 1, 6, 16] and error detection models [3, 5, 8]. However, existing mistake detection approaches vary widely. Some methods emphasize action detection, identifying specific errors like missing steps, out-of-sequence actions [17] or incorrect ordering [16], while other models bypass action detection altogether and monitor changes to the assembled object to ensure procedural accuracy [18, 19]. This variability in proposed approaches and their intended objectives highlights the need for more comprehensive evaluation frameworks and standardized benchmarks to assess and compare the effectiveness of different mistake detection strategies and for more olistic models, able to detect any deviation from the correct workflow.
@@
An ideal Procedural Mistake Detection (PMD) model should demonstrate two critical capabilities: robustness to diverse mistake types and the ability to provide online, timely feedback. Regarding robustness, a reliable PMD system should be capable of identifying any mistake that may occur within a structured procedure, regardless of its type. To uphold this principle, in this paper we propose framing mistake detection as a One-Class Classification problem, where the model is trained to learn the correct sequences of steps within a procedure and is then able to recognize any deviation from the correct workflow. Within the OCC framework, during training, the PMD model is only exposed to “normal” procedures [20, 21] to learn the correct workflows, i.e. the sequences of performed steps that allow to correctly complete the task. Conversely, during inference, the learned model should detect any deviation from the learnt workflow, deeming that step as an error. Further, it is worth noting that the OCC framework requires only binary labels (Correct vs. Mistake), which are easier to collect with respect to fine-grained labels like the error type when labelling videos of procedural datasets. Second, the model should provide online [^1] We distinguish an online setup, which allows slight processing delays while enabling effective feedback, from a real-time setup, which requires immediate responses. feedback quickly enough to allow users to correct mistakes promptly, minimizing the reinforcement of incorrect actions and reducing the risk of hazardous situations. Besides these two pivotal abilities, the PMD system should leverage an egocentric perspective in procedural videos, enhancing the model’s applicability by mirroring natural human perception, and allowing for seamless real-world deployment on wearable devices.
@@
This paper extends our previous conference work, PREGO [22], by refining key components and methods to improve procedural mistake detection (see Fig. 1). Unlike our original study, which primarily reported classical Precision and Recall metrics, in this work, we adopt Balanced Precision [23] to mitigate the extreme class imbalance inherent in egocentric mistake detection. Balanced Precision scales the False Positives by the ratio of positive‐to‐negative samples, ensuring that the penalty for mislabeling a correct step is placed on the same scale as the reward for correctly identifying a mistake. Specifically, we lead a comprehensive evaluation of several LLMs as step anticipators, increasing the state‑of‑the‑art results on Assembly101‑O and Epic Tent‑O [22]. We compare a broad range of LLMs such as Llama‑2 [24], Llama‑3 [25], Gemma [26], and Mistral [27], fine‑tune them using Low‑Rank Adaptation (LoRA), and experiment with prompting methods including Zero‑Shot, Few‑Shot, and Auto‑CoT [28].
@@ 2.1 Procedural Mistake Detection
Procedural learning has seen notable progress due to the development of diverse datasets [6, 10, 9, 11, 12], which offer valuable insights into both structured [7, 3, 29] and unstructured tasks [1, 30]. These datasets span a wide range of applications—from industrial assembly [29, 2, 8, 12] to everyday cooking activities [30, 31, 32]—yet the lack of a standardized benchmark for mistake detection has led to limited evaluation and a fragmented literature.
In Table 1, we compare relevant models for procedural mistake detection across multiple dimensions: whether they operate in an open-set or one-class condition, whether they support online or offline detection, the data modalities they leverage (e.g., RGB frames, hand poses, eye gaze, keystep logic), the task they address, and the datasets used. Notably, models marked with an asterisk (*) follow our protocol introduced in [22], standardizing an online, open-set detection scenario for consistent evaluation.
Existing methods illustrate a variety of approaches. For instance, Ding et al. [5] employ knowledge graphs with textual transcripts, neglecting visual cues. Ghoddoosian et al. [4] focus on action recognition, treating error detection as a semantic evaluation of segmentation results. Other works, such as Assembly101 [2] and HoloAssist [3], apply error detection baselines offline, relying on pre-segmented videos and annotated labels. By contrast, [8] emphasizes Procedure Step Recognition, focusing on correct step completion rather than partial activities, while EgoPER [33] integrates action segmentation with contrastive learning to detect previously unseen errors. Our prior work, PREGO [22], uses RGB frames alongside symbolic reasoning to assess procedural correctness online, operating under an OCC framework and trained solely on error-free sequences [20, 21]. Finally, Seminara et al. [16] adopt a distinct approach, learning task graphs via Maximum Likelihood estimation of key-step sequences.
@@ 2.2 Datasets of Procedural Egocentric Videos
Previous investigations proposed datasets and benchmarks to support research in procedural mistake detection. IndustReal [8] focuses on a single toy, resulting in the learning of only one procedure without annotations for procedural mistakes. Assembly101 [2] is a large-scale video dataset that provides frame-level mistake annotations. It features videos of actors assembling toys, with synchronized Ego-Exo views and annotated hand positions. IndustReal [8] focuses on a single toy, resulting in the learning of only one procedure without annotations for procedural mistakes. Epic-tent [1] is a dataset that covers a different domain of unscripted actions that capture actors assembling a tent outdoors. The participants exhibit varying levels of expertise and naturally make mistakes, which have been annotated. ATA [4] is a procedural dataset created for offline mistake detection in assembling activities. However, it provides only video-level mistake annotations, limiting its usefulness for frame-based applications. HoloAssist [3] provides egocentric videos of individuals performing multiple manipulation tasks following expert instructions. CaptainCook4D [34] is tailored for evaluating procedural activities, specifically focusing on the culinary domain.
Notably,  [33] introduced a novel egocentric procedural error dataset consisting of videos depicting various errors within the cooking domain.
In this study, we propose a new benchmark building upon the datasets from [2, 1] as they provide insights into procedural errors in two distinct settings: a controlled industrial environment and outdoor environments.
Table 1: Comparison of relevant models in procedural mistake detection. In the modalities column, RGB refers to RGB images, H stands for hand poses, E represents eye gaze and K indicates keystep labels. Our approach is the first to adopt an egocentric, one-class and online method for detecting procedural mistakes. (*) denotes that these models are based on our protocol defined in [22].
@@
Ding et al. [5] - ArXiv ’23 K Mistake Detection Assembly101 [2]
Wang et al. [3] - ICCV ’23 RGB+H+E Mistake Detection HoloAssist [3]
Ghoddoosian et al. [4] - ICCV ’23 RGB Unknown Sequence Detection ATA [4], CSV [35]
Schoonbeek et al. [8] - WACV ’24 Multi Procedure Step Recognition IndustReal [8]
Lee et al. [33] - CVPR ’24 RGB Mistake Detection EgoPER [33], HoloAssist [3], ATA [4]
Seminara et al. [16] - NeurIPS’24 * RGB Mistake Detection Assembly101-O, Epic-tent-O
PREGO - CVPR ’24 * [22] RGB Mistake Detection Assembly101-O, Epic-tent-O
@@ 2.3 Step Recognition
Zhuang et al. [13] introduce an action segmentation model that employs an attention-based architecture coupled with a Pairwise Ordering Consistency (POC) loss function. Moreover, they develop a weakly-supervised method that requires only the set of actions in a procedure as input, eliminating the need for labor-intensive frame-level annotations. In Shah et al. [36], a novel loss for self-supervised learning is combined with a clustering algorithm to detect key steps in procedural videos without using any labels. [15] tackles the task by utilizing online instructional videos to learn actions and sequences without manual annotations, combining step recognition with a deep probabilistic model to account for step order and timing variability.
An et al. [37] introduced MiniROAD, the state-of-the-art online step recognizer, which leverages an RNN architecture and adjusts loss importance during training to perform online action recognition. The approach proposed in Shen et al. [38] also highlights the importance of holistically understanding actions rather than treating each frame in isolation. In contrast to previous methods that rely on frame-by-frame detection—such as MiniROAD [37], which focuses on step recognition—our approach addresses the limitations of isolated frame-level analysis. While prior work has established a strong foundation in procedural step recognition, it often falls short in capturing the broader temporal context that is critical in real-world scenarios. To overcome this, we propose and evaluate several aggregation techniques that enhance the coherence of action recognition across sequences, thereby enabling more robust and context-aware mistake detection.
@@ 2.4 Step Anticipation
[39] generates multiple possible natural language outcomes, leveraging pretraining on a text corpus to deal with the variability in future actions. [40] employs a two-level hierarchical approach, integrating both high-level human intentions and low-level action sequences, improving action anticipation in the long term. The pipeline consists of two models: the first one extracts intentions and classifies actions while the second one generates future actions, conditioning human intentions to narrow down the uncertainty set of future actions. The framework presented in [41] addresses future activity anticipation in egocentric videos by employing a contrastive loss to emphasize novel information and a dynamic reweighting mechanism that prioritizes informative past content. This approach improves video representation, leading to more accurate predictions of future activities.
Similar to AntGPT [42], we leverage video frames to anticipate future actions using an LLM enhanced with Automatic Chain of Thought. However, while previous approaches, such as AntGPT, require predefined segmentation of a long video into ordered, annotated short segments with corresponding action labels, our online setting processes each frame independently without prior segmentation.
@@ 2.5 Reasoning Tasks with Large Language Models
Large Language Models (LLMs), trained on vast datasets and equipped with numerous parameters, exhibit advanced capabilities beyond those of earlier language models [43]. They have demonstrated exceptional performance in both natural language processing tasks [24] and non-language tasks [44, 43, 45]. The next-token prediction mechanism of LLMs closely parallels our action anticipation framework, as both aim to predict future outcomes based on accumulated data.
Previous research has shown that LLMs can generate semantically meaningful patterns [46, 42], while other work [47] has explored their in-context learning abilities with semantically unrelated labels, where there is no direct relationship between a token and its meaning. More recent work  [48, 45, 49, 50, 46, 51, 52] has further investigated the ability of LLMs to function as In-Context Learners (ICLs), meaning they can solve novel and unseen tasks relying only on a structured and informative prompt, without requiring additional fine-tuning. When provided with a query prompt that includes a context of input-output examples, LLMs can understand the problem and generate appropriate solutions within this framework. LLMs as ICLs have been applied to a wide range of tasks, including planning [48, 52], programming [45, 49], logical problem-solving [50] and symbolic reasoning [46]. In addition to ICL, other paradigms, such as Chain-of-Thought (CoT) [53] and Automatic Chain-of-Thought (ACoT) [28], have been employed to enable LLMs to reason through their responses step-by-step, improving their ability to plan and articulate their thought process explicitly.
Palm [51] is the most similar to our anticipation module, as it employs a Socratic method [54] that uses video features to predict subsequent actions with an LLM. Socratic Models are a framework that utilizes structured dialogue between pre-existing foundation models, each leveraging its unique capabilities based on its training data distribution. However, while PALM  [55] integrates an Action Recognition Model and a Vision-Language Model to generate text prompts for long-term action anticipation, we diverge by using a dual-branch architecture: one branch tasked to recognize the current action and the other leveraging an LLM for action anticipation focusing on next step prediction. This schema offer a natural pipeline for error detection by comparing the predictions generated by the two branches

### v3

@@ 1 Introduction
Figure 1: Procedural Mistake Detection involves identifying errors within a procedural video. Each procedure is composed of different steps that should be executed with a certain order. The aim is to develop a method capable of analyzing a video and determining whether each frame contains a mistake, e.g., a step that is not executed in the correct order. Addressing this task is particularly useful when applied to wearable devices, as they allow for direct feedback to the person performing the task. In the Figure, TI-PREGO, depicted by a gear icon, takes as input the video sequence from time 0 to t1t-1, classifying the current frame $t$ as either correct (green) or a mistake (red). This process continues frame by frame until a mistake is detected.
@@
Advanced mistake detection models will become essential to ensure precision, safety, and efficiency in procedural applications. This demand has recently resulted in an increasing number of datasets [epicTent, sener2022assembly101, HoloAssist2023, Ghoddoosian_2023_ICCV, ding2023every, coin, ikea, IndustReal, cross_task, proceL_2019, howto100m_2019_miech, ragusa2023enigma51] and methodologies designed to advance procedure learning [POC_2022, Ghoddoosian_2023_ICCV, OadTR, zhong_2023, epicTent, coin, seminara2024differentiable] and error detection models [HoloAssist2023, ding2023every, IndustReal]. However, existing mistake detection approaches vary widely. Some methods emphasize action detection, identifying specific errors like missing steps, out-of-sequence actions [narasimhan2023learning] or incorrect ordering [seminara2024differentiable], while other models bypass action detection altogether and monitor changes to the assembled object to ensure procedural accuracy [chen2020monitoring, candido2023image]. This variability in proposed approaches and their intended objectives highlights the need for more comprehensive evaluation frameworks and standardized benchmarks to assess and compare the effectiveness of different mistake detection strategies and for more olistic models, able to detect any deviation from the correct workflow.
@@
An ideal Procedural Mistake Detection (PMD) model should demonstrate two critical capabilities: robustness to diverse mistake types and the ability to provide online, timely feedback. Regarding robustness, a reliable PMD system should be capable of identifying any mistake that may occur within a structured procedure, regardless of its type. To uphold this principle, in this paper we propose framing mistake detection as a One-Class Classification (OCC) problem, where the model is trained to learn the correct sequences of steps within a procedure and is then able to recognize any deviation from the correct workflow. Within the OCC framework, during training, the PMD model is only exposed to “normal” procedures [Flaborea_2023_ICCV, zaheer2022generative] to learn the correct workflows, i.e. the sequences of performed steps that allow to correctly complete the task. Conversely, during inference, the learned model should detect any deviation from the learnt workflow, deeming that step as an error. Further, it is worth noting that the OCC framework requires only binary labels (Correct vs. Mistake), which are easier to collect with respect to fine-grained labels, such as the error type, when labelling videos of procedural datasets. Second, the model should provide online [^1] We distinguish an online setup, which allows slight processing delays while enabling effective feedback, from a real-time setup, which requires immediate responses. feedback quickly enough to allow users to correct mistakes promptly, minimizing the reinforcement of incorrect actions and reducing the risk of hazardous situations. It is worth noting that the requirement for immediate feedback defines the online setting and imposes a strict causal constraint on the PMD models as they are allowed to leverage only past and present information for producing their decisions. In contrast, previous methods [Ghoddoosian_2023_ICCV, ding2023every, HoloAssist2023, lee2024error] adopt an offline framework, where the analysis is performed using complete pre-recorded procedures. This allows to employ bidirectional models that leverage both past and future context, achieving high accuracy in retrospective error identification for post-hoc analysis like performance grading, but making them unsuitable for live intervention. Thus, we argue that the processing approach must shift towards proactive error anticipation, where the model predicts the next correct step of the sequence to timely flag deviations. Besides these two pivotal abilities, the PMD system should leverage an egocentric perspective in procedural videos, enhancing the model’s applicability by mirroring natural human perception, and allowing for seamless real-world deployment on wearable devices.
@@
This paper extends our previous conference work, PREGO [Flaborea_2024_CVPR], by refining key components and methods to improve procedural mistake detection (see Fig. 1). Unlike our original study, which primarily reported classical Precision and Recall metrics, in this work, we adopt Balanced Precision [Wang2021OadTROA] to mitigate the extreme class imbalance inherent in egocentric mistake detection. Balanced Precision scales the False Positives by the ratio of positive‐to‐negative samples, ensuring that the penalty for mislabeling a correct step is placed on the same scale as the reward for correctly identifying a mistake. Specifically, we lead a comprehensive evaluation of several LLMs as step anticipators, increasing the state‑of‑the‑art results on Assembly101‑O and Epic Tent‑O [Flaborea_2024_CVPR]. We compare a broad range of LLMs such as Llama‑2 [touvron2023llama], Llama‑3 [llama3], Gemma [gemma], and Mistral [mistral7B], fine‑tune them using Low‑Rank Adaptation (LoRA), and experiment with prompting methods including Zero‑Shot, Few‑Shot, and Auto‑CoT [Zhang2022AutomaticCO].
@@ 2.1 Procedural Mistake Detection
Procedural learning has seen notable progress due to the development of diverse datasets [coin, proceL_2019, cross_task, howto100m_2019_miech, ragusa2023enigma51], which offer valuable insights into both structured [ikea, HoloAssist2023, Ragusa_meccano_2021] and unstructured tasks [epicTent, epickitchens_Damen_2018]. These datasets span a wide range of applications—from industrial assembly [Ragusa_meccano_2021, sener2022assembly101, IndustReal, ragusa2023enigma51] to everyday cooking activities [epickitchens_Damen_2018, breakfasts, salads]—yet the lack of a standardized benchmark for mistake detection has led to limited evaluation and a fragmented literature.
In Table 1, we compare relevant models for procedural mistake detection across multiple dimensions: whether they operate in an open-set or one-class condition, whether they support online or offline detection, the data modalities they leverage (e.g., RGB frames, hand poses, eye gaze, keystep logic), the task they address, and the datasets used. Notably, models marked with an asterisk (*) follow our protocol introduced in [Flaborea_2024_CVPR], standardizing an online, open-set detection scenario for consistent evaluation.
Existing methods illustrate a variety of approaches. For instance, Ding et al. [ding2023every] employ knowledge graphs with textual transcripts, neglecting visual cues. Ghoddoosian et al. [Ghoddoosian_2023_ICCV] focus on action recognition, treating error detection as a semantic evaluation of segmentation results. Other works, such as Assembly101 [sener2022assembly101] and HoloAssist [HoloAssist2023], apply error detection baselines offline, relying on pre-segmented videos and annotated labels. By contrast, [IndustReal] emphasizes Procedure Step Recognition, focusing on correct step completion rather than partial activities, while EgoPER [lee2024error] integrates action segmentation with contrastive learning to detect previously unseen errors. Our prior work, PREGO [Flaborea_2024_CVPR], uses RGB frames alongside symbolic reasoning to assess procedural correctness online, operating under an OCC framework and trained solely on error-free sequences [Flaborea_2023_ICCV, zaheer2022generative]. Finally, Seminara et al. [seminara2024differentiable] adopt a distinct approach, learning task graphs via Maximum Likelihood estimation of key-step sequences.
@@ 2.2 Datasets of Procedural Egocentric Videos
Previous investigations proposed datasets and benchmarks to support research in procedural mistake detection. IndustReal [IndustReal] focuses on a single toy, resulting in the learning of only one procedure without annotations for procedural mistakes. Assembly101 [sener2022assembly101] is a large-scale video dataset that provides frame-level mistake annotations. It features videos of actors assembling toys, with synchronized Ego-Exo views and annotated hand positions. IndustReal [IndustReal] focuses on a single toy, resulting in the learning of only one procedure without annotations for procedural mistakes. Epic-tent [epicTent] is a dataset that covers a different domain of unscripted actions that capture actors assembling a tent outdoors. The participants exhibit varying levels of expertise and naturally make mistakes, which have been annotated. ATA [Ghoddoosian_2023_ICCV] is a procedural dataset created for offline mistake detection in assembling activities. However, it provides only video-level mistake annotations, limiting its usefulness for frame-based applications. HoloAssist [HoloAssist2023] provides egocentric videos of individuals performing multiple manipulation tasks following expert instructions. CaptainCook4D [peddi2023captaincook4d] is tailored for evaluating procedural activities, specifically focusing on the culinary domain. Notably,  [lee2024error] introduced a novel egocentric procedural error dataset consisting of videos depicting various errors within the cooking domain.
In this study, we propose a new benchmark building upon the datasets from [sener2022assembly101, epicTent] as they provide insights into procedural errors in two distinct settings: a controlled industrial environment and outdoor environments.
Table 1: Comparison of relevant models in procedural mistake detection. In the modalities column, RGB refers to RGB images, H stands for hand poses, E represents eye gaze and K indicates keystep labels. Our approach is the first to adopt an egocentric, one-class and online method for detecting procedural mistakes. (*) denotes that these models are based on our protocol defined in [Flaborea_2024_CVPR].
@@
Ding et al. [ding2023every] - ArXiv ’23 K Mistake Detection Assembly101 [sener2022assembly101]
Wang et al. [HoloAssist2023] - ICCV ’23 RGB+H+E Mistake Detection HoloAssist [HoloAssist2023]
Ghoddoosian et al. [Ghoddoosian_2023_ICCV] - ICCV ’23 RGB Unknown Sequence Detection ATA [Ghoddoosian_2023_ICCV], CSV [Qian_2022_CVPR]
Schoonbeek et al. [IndustReal] - WACV ’24 Multi Procedure Step Recognition IndustReal [IndustReal]
Lee et al. [lee2024error] - CVPR ’24 RGB Mistake Detection EgoPER [lee2024error], HoloAssist [HoloAssist2023], ATA [Ghoddoosian_2023_ICCV]
Seminara et al. [seminara2024differentiable] - NeurIPS’24 * RGB Mistake Detection Assembly101-O, Epic-tent-O
PREGO - CVPR ’24 * [Flaborea_2024_CVPR] RGB Mistake Detection Assembly101-O, Epic-tent-O
@@ 2.3 Step Recognition
Zhuang et al. [POC_2022] introduce an action segmentation model that employs an attention-based architecture coupled with a Pairwise Ordering Consistency (POC) loss function. Moreover, they develop a weakly-supervised method that requires only the set of actions in a procedure as input, eliminating the need for labor-intensive frame-level annotations. In Shah et al. [Shah_2023_STEPs], a novel loss for self-supervised learning is combined with a clustering algorithm to detect key steps in procedural videos without using any labels. [zhong_2023] tackles the task by utilizing online instructional videos to learn actions and sequences without manual annotations, combining step recognition with a deep probabilistic model to account for step order and timing variability.
An et al. [An2023MiniROADMR] introduced MiniROAD, the state-of-the-art online step recognizer, which leverages an RNN architecture and adjusts loss importance during training to perform online action recognition. The approach proposed in Shen et al. [shen2024progress] also highlights the importance of holistically understanding actions rather than treating each frame in isolation. In contrast to previous methods that rely on frame-by-frame detection—such as MiniROAD [An2023MiniROADMR], which focuses on step recognition—our approach addresses the limitations of isolated frame-level analysis. While prior work has established a strong foundation in procedural step recognition, it often falls short in capturing the broader temporal context that is critical in real-world scenarios. To overcome this, we propose and evaluate several aggregation techniques that enhance the coherence of action recognition across sequences, thereby enabling more robust and context-aware mistake detection.
@@ 2.4 Step Anticipation
[Abdelsalam_2023_gepsan] generates multiple possible natural language outcomes, leveraging pretraining on a text corpus to deal with the variability in future actions. [mascaro2023intention] employs a two-level hierarchical approach, integrating both high-level human intentions and low-level action sequences, improving action anticipation in the long term. The pipeline consists of two models: the first one extracts intentions and classifies actions while the second one generates future actions, conditioning human intentions to narrow down the uncertainty set of future actions. The framework presented in [self_regulated] addresses future activity anticipation in egocentric videos by employing a contrastive loss to emphasize novel information and a dynamic reweighting mechanism that prioritizes informative past content. This approach improves video representation, leading to more accurate predictions of future activities.
Similar to AntGPT [Zhao2023AntGPTCL], we leverage video frames to anticipate future actions using an LLM enhanced with Automatic Chain of Thought. However, while previous approaches, such as AntGPT, require predefined segmentation of a long video into ordered, annotated short segments with corresponding action labels, our online setting processes each frame independently without prior segmentation.
@@ 2.5 Reasoning Tasks with Large Language Models
Large Language Models (LLMs), trained on vast datasets and equipped with numerous parameters, exhibit advanced capabilities beyond those of earlier language models [Wei2022EmergentAO]. They have demonstrated exceptional performance in both natural language processing tasks [touvron2023llama] and non-language tasks [brooks2023instructpix2pix, Wei2022EmergentAO, gupta2023visual]. The next-token prediction mechanism of LLMs closely parallels our action anticipation framework, as both aim to predict future outcomes based on accumulated data.
Previous research has shown that LLMs can generate semantically meaningful patterns [generalpatternmachines2023, Zhao2023AntGPTCL], while other work [Wei2023LargerLM] has explored their in-context learning abilities with semantically unrelated labels, where there is no direct relationship between a token and its meaning. More recent work  [Pallagani2022PlansformerGS, gupta2023visual, Liang2022CodeAP, Feng2023LanguageMC, generalpatternmachines2023, Kim2024PALMPA, Ahn2022DoAI] has further investigated the ability of LLMs to function as In-Context Learners (ICLs), meaning they can solve novel and unseen tasks relying only on a structured and informative prompt, without requiring additional fine-tuning. When provided with a query prompt that includes a context of input-output examples, LLMs can understand the problem and generate appropriate solutions within this framework. LLMs as ICLs have been applied to a wide range of tasks, including planning [Pallagani2022PlansformerGS, Ahn2022DoAI], programming [gupta2023visual, Liang2022CodeAP], logical problem-solving [Feng2023LanguageMC] and symbolic reasoning [generalpatternmachines2023]. In addition to ICL, other paradigms, such as Chain-of-Thought (CoT) [Wei2022ChainOT] and Automatic Chain-of-Thought (ACoT) [Zhang2022AutomaticCO], have been employed to enable LLMs to reason through their responses step-by-step, improving their ability to plan and articulate their thought process explicitly.
Palm [Kim2024PALMPA] is the most similar to our anticipation module, as it employs a Socratic method [Zeng2022SocraticMC] that uses video features to predict subsequent actions with an LLM. Socratic Models are a framework that utilizes structured dialogue between pre-existing foundation models, each leveraging its unique capabilities based on its training data distribution. However, while PALM  [chowdhery2023palm] integrates an Action Recognition Model and a Vision-Language Model to generate text prompts for long-term action anticipation, we diverge by using a dual-branch architecture: one branch tasked to recognize the current action and the other leveraging an LLM for action anticipation focusing on next step prediction. This schema offer a natural pipeline for error detection by comparing the predictions generated by the two branches